OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! What are the limits of integration if the summation the limit as n goes to infinity of the summation from k equals 1 to n of the product of the quantity of the square of 2 plus 7 times k over n and the quotient of 7 and n is written as a definite integral with integrand x2? a. 0,2 b. 1,7 c. 2,9 d. 2,7

OpenStudy (xguardians):

Can you post the equation out, it makes it easier to understand.

OpenStudy (aryana_maria2323):

\[\lim_{n \rightarrow \infty}\sum_{k=1}^{n}[2+\frac{ 7k }{ n }]^2(\frac{ 7 }{ n })\]

OpenStudy (xguardians):

http://openstudy.com/study#/updates/572512e6e4b0644381e0fd45 Reference this thread.

OpenStudy (xguardians):

We solved something very similar and it has lots of resources.

OpenStudy (aryana_maria2323):

I don't understand.

OpenStudy (aryana_maria2323):

@xGuardians Can you show me how to solve this in steps?

OpenStudy (xguardians):

Quite honestly, I just learned this stuff today :P

OpenStudy (xguardians):

\[\int\limits_{?}^{?}(2+x)^2\]

OpenStudy (aryana_maria2323):

Okay then you should probably remember it very well lol

OpenStudy (xguardians):

Well, I didn't really learn it, I learned it in that thread I referenced you to...

OpenStudy (aryana_maria2323):

Can you try then?

OpenStudy (xguardians):

Oh, I get it.

OpenStudy (xguardians):

So you see the 7/n part, that means the limits of integration must be 7 apart.

OpenStudy (xguardians):

0,2 = 2 apart 1,7 = 6 apart 2,9 = 7 apart <---- 2,7 = 5 apart

OpenStudy (aryana_maria2323):

okay so 2,9 is my answer?

OpenStudy (xguardians):

Yeppers.